OpenStudy (anonymous):

Simplify. Please show work. (-4x)^4

OpenStudy (anonymous):

do it one step at a time -4^4=256 and x^4=x^4 bring them together to get 256x^4

OpenStudy (lilg132):

-4 ^ 4 = 256 so it is 256x^4 Please remember to give a medal

OpenStudy (anonymous):

Of course ill give you a medal:) You are the best.